# A Hybrid Jump–Diffusion Model for Coherent Optical Control of Quantum Emitters in hBN

### v1

@@ Abstract
Hexagonal boron nitride (hBN) has emerged as a promising two-dimensional host for stable single-photon emission owing to its wide bandgap, high photostability, and compatibility with nanophotonic integration. We present a simulation-based study of temperature-dependent spectral dynamics and optical coherence in a mechanically decoupled quantum emitter in hBN. Employing a hybrid stochastic framework that combines Ornstein–Uhlenbeck detuning fluctuations with temperature-dependent, Gaussian-distributed discrete frequency jumps, motivated by experimentally observed spectral diffusion and blinking, we reproduce the measured evolution of inhomogeneous linewidth broadening and the progressive degradation of photon coherence across the relevant cryogenic range (5–30 K). The model captures phonon-related spectral diffusion with a cubic temperature dependence and the onset of jump-like spectral instabilities at higher temperatures. By calibrating the hybrid diffusion, jump parameters to the experimentally measured full width at half maximum (FWHM) of the emission line and analyzing the second-order correlation function $g^{(2)}(\tau)$ under resonant driving, we establish a unified phenomenological description that links stochastic detuning dynamics to the decay of optical coherence in a resonantly driven emitter. Analysis of $g^{(2)}(\tau)$ under resonant driving reveals an additional dephasing rate $\gamma_{\mathrm{sd+j}}$ that rises monotonically with temperature and drive strength, leading to a predicted critical crossover to overdamped dynamics at $T_{\mathrm{crit}}\approx 25.91$ K. This hybrid framework provides a quantitative connection between accessible spectroscopic observables and the dominant noise mechanisms limiting coherent optical control in mechanically decoupled quantum emitters, exemplified in hBN and generalizable to similar emitters in other materials.
@@ I Introduction
hBN possesses a wide bandgap of approximately $5.9~\text{eV}$ [1, 2], supporting deep-ultraviolet to near-infrared emission arising from defect-associated electronic transitions [3, 4, 5]. Owing to their brightness, polarization stability, and integrability, hBN quantum emitters have emerged as promising alternatives to semiconductor quantum dots for stable single-photon sources, provided that appropriate device engineering (encapsulation, substrate choice, or mechanical decoupling) is applied to mitigate spectral diffusion and sample variability [6, 7, 8]. In mechanically decoupled quantum emitters in hBN, reduced coupling to substrate-induced strain and charge noise has been shown to result in enhanced spectral stability and narrow optical resonances at cryogenic temperatures, including near 660 nm in representative systems [1]. This mechanical isolation enables coherent optical control and exposes intrinsic phonon-limited coherence that is otherwise masked in substrate-bound samples. Despite these advantages, the optical coherence of hBN defect centers remains strongly influenced by phononic and electrostatic fluctuations. Temperature-dependent spectroscopy has shown that spectral diffusion, phonon coupling, and charge-induced dephasing can jointly broaden the emission linewidth and degrade coherent control of emitters in hBN [10, 11, 12]. For specific classes of emitters [1], both the zero-phonon line (ZPL) shift and linewidth have been reported to follow a cubic temperature dependence, consistent with deformation-potential coupling to low-energy acoustic phonons [13]. However, alternative temperature scalings have been reported for other hBN emitters, reflecting defect-specific phonon-coupling mechanisms and local environments [14]. Photoluminescence excitation spectroscopy further reveals that, in such emitters, an energy gap between the ZPL and the first acoustic mode of the phonon sideband can close with increasing temperature, leading to a rapid increase in dephasing once phonon protection is lost [1]. Beyond homogeneous phonon-induced broadening, hBN emitters also exhibit significant inhomogeneous broadening caused by slow spectral diffusion. Local charge reconfigurations, strain fluctuations, and defect rearrangements induce stochastic frequency shifts that broaden time-averaged spectra beyond the lifetime-limited linewidth, even at cryogenic temperatures. For the mechanically decoupled quantum emitter in hBN that serves as the experimental reference throughout this work (Ref. [1]), the measured linewidth follows
@@
with $A=1.01~\mathrm{GHz}$ and $B=3.77\times 10^{-5}~\mathrm{GHz/K}^{3}$. The constant term reflects radiative lifetime broadening together with residual quasi-static spectral diffusion, while the cubic term arises from acoustic-phonon-mediated dephasing. Mechanical decoupling suppresses the temperature-independent inhomogeneous component, but at higher temperatures, thermally activated charge motion and lattice fluctuations reintroduce spectral wandering and discrete frequency jumps that dominate the emitter’s dynamics (see Supplementary Information, Sec. III). Coherent optical control—including Rabi oscillations and $\pi$-pulse excitation—has been demonstrated up to $\sim 30~\mathrm{K}$ in mechanically isolated hBN [1, 15, 16]. At elevated temperatures, however, activated lattice noise and charge fluctuations lead to a transition from coherent to overdamped behavior governed by both continuous spectral wandering and abrupt jump-like instabilities [17, 18]. Reports of near-Fourier-limited coherence approaching room temperature in certain hBN-based emitter platforms, including suspended and nanophotonic device geometries, highlight the strong sensitivity of coherence preservation to fabrication quality, suspension geometry, and the local nano-environment [10, 19, 16, 1]. In this work, we develop a hybrid stochastic model inspired by recent theoretical treatments of spectral diffusion in solid-state emitters [3, 2], combining Ornstein–Uhlenbeck diffusion with a discrete Gaussian jump process to capture both continuous and abrupt detuning fluctuations (see Supplementary Information, Sec. I for numerical implementation and Sec. III for the physical motivation of jumps). Using experimentally extracted temperature-dependent parameters, the model reproduces the measured linewidth evolution, the power-dependent damping of optical Rabi oscillations, and the temperature-dependent behavior of the second-order correlation function $g^{(2)}(\tau)$ (see Supplementary Information, Fig. S2 and Sec. VII). This unified framework links phonon-driven diffusion with jump-like spectral instabilities and enables quantitative simulation of coherence dynamics in mechanically decoupled hBN emitters.
@@ II Theoretical Model and Simulation Framework
The spectral dynamics of a single quantum emitter in hBN are modeled using a hybrid stochastic approach that captures both continuous and discrete detuning fluctuations. This framework is motivated by recent theoretical treatments of spectral diffusion in solid-state emitters [3, 2] and by foundational studies of Gaussian and telegraph-like noise processes in semiconductor defects. The model combines Ornstein–Uhlenbeck (OU) diffusion, representing phonon-induced dephasing, with a Gaussian Random Jump (GRJ) process describing abrupt frequency shifts associated with charge motion, defect rearrangements, or local strain reconfigurations (see Supplementary Information, Sec. III). In the discrete form implemented numerically, the instantaneous detuning $\omega_{i}$ evolves according to
@@
where $dW_{t}$ is the Wiener increment and $J(t)$ is a compound Poisson process with rate $\lambda_{J}$ and Gaussian jump amplitudes. This hybrid diffusion–jump dynamics produces an effective inhomogeneously broadened emission line whose full width at half maximum (FWHM) is approximated by
@@
as measured in the mechanically decoupled hBN emitter of Koch et al. [1] (see Supplementary Information, Sec. VI and Table I for the complete two-stage calibration protocol using physically motivated temperature-dependent multipliers). The parameters $(S,\lambda_{J},\sigma_{J})$ are treated as temperature dependent and increase monotonically with $T$. For each temperature, Eq. (2) is integrated numerically using the Euler–Maruyama method over a 10 ns window, and the resulting detuning histograms are fitted to extract $\Gamma_{\mathrm{sim}}(T)$ (see Supplementary Information, Sec. V).
@@ III Simulation Results
The calibrated jump–diffusion model reproduces the experimentally observed temperature dependence of the emission linewidth of the specific mechanically decoupled hBN quantum emitter studied in Ref. [1]. For each temperature, the stochastic parameters $(S,\lambda_{J},\sigma_{J})$ were obtained through a constrained calibration in which their values were adjusted monotonically to match the experimentally measured inhomogeneous broadening $\Gamma_{\mathrm{exp}}(T)=A+BT^{3}$ with an accuracy better than $0.02~\mathrm{GHz}$ (see Supplementary Information, Sec. VI). The calibrated parameters provide qualitative insight into the relative contributions of continuous diffusion and jump-like detuning processes, and the resulting agreement between simulation and experiment is shown in Fig. 1, where the simulated linewidths reproduce the cubic, phonon-limited trend across the experimentally relevant temperature range from 5 to 30 K.
@@
Figure 2: (a–c) Simulated second-order correlation functions $g^{(2)}(\tau)$ for the hBN quantum emitter studied in Ref. [1] at 5 K, 20 K, and 30 K, respectively. Increasing excitation power increases the Rabi frequency ($\Omega_{R}\propto\sqrt{P}$), which manifests in $g^{(2)}(\tau)$ as more pronounced oscillatory shoulders and a higher oscillation frequency at short delay times. At elevated temperatures, these oscillatory features are progressively damped due to enhanced spectral diffusion and thermally activated frequency jumps, leading to suppression of coherent Rabi dynamics (see Supplementary Information, Fig. S2 for parameter dependence).
The effective decay rate extracted from the simulated photon-correlation traces provides a direct quantitative measure of how spectral diffusion and jump processes modify coherent optical dynamics. The additional dephasing term $\gamma_{\mathrm{sd+j}}$ appearing in Eq. (6) encapsulates the combined influence of continuous OU-type spectral diffusion and discrete Gaussian jumps—precisely the same mechanisms responsible for the inhomogeneous broadening discussed earlier in the linewidth analysis. As shown in Fig. 3, $\gamma_{\mathrm{sd+j}}$ is negligible at 5 K, resulting in decay rates that remain nearly independent of the Rabi frequency. At 20 K, $\gamma_{\mathrm{sd+j}}$ increases from a negligible value and develops a weak but measurable linear dependence on the Rabi frequency $\Omega_{R}$, coinciding with a fivefold increase in the calibrated jump amplitude $\sigma_{J}$, while the diffusion strength $S$ and jump rate $\lambda_{J}$ remain unchanged. At 30 K, $\gamma_{\mathrm{sd+j}}$ increases further as the jump-induced detuning remains strong and the diffusion strength $S$ rises by approximately 10%, leading to a steep increase in the decay rate that surpasses the $\Gamma=\Omega_{R}$ boundary. This marks the transition to an overdamped regime ($\Gamma=2\times\Omega_{R}$), in which noise-induced detuning dominates over coherent Rabi cycling. The resulting monotonic, temperature-driven enhancement of $\gamma_{\mathrm{sd+j}}$ establishes a direct link between inhomogeneous broadening, set by the combined action of spectral diffusion and jump processes. The observed suppression of coherent dynamics under resonant driving (see Supplementary Information, Fig. S3, which demonstrates that only the hybrid model with discrete jumps reproduces the experimentally observed compact, near-Gaussian detuning distributions at elevated temperatures).
@@ IV Conclusion
Our hybrid jump-diffusion model reproduces the experimentally observed temperature and power-dependent coherence degradation of the mechanically decoupled hBN quantum emitter studied in Ref. [1], using only the measured $T^{3}$ inhomogeneous broadening law as input [1]. The framework captures the interplay between continuous spectral diffusion and discrete jump-like detuning processes responsible for coherence loss under resonant driving. While the numerical values of the calibrated parameters are specific to this emitter, the modeling approach can be extended to mechanically decoupled hBN quantum emitters exhibiting similar temperature-dependent inhomogeneous broadening. Full numerical implementation, parameter scans, calibration strategy, analytical derivations, and validation are provided in the Supplementary Information (Secs. I–VII, Eqs. (1)–(9), Figs. S1–S3, and Table I). Inversion of Calibration and Extraction of Critical Dephasing Temperature
Most importantly, the model identifies the fundamental crossover where the effective dephasing rate equals the Rabi frequency (slope = 1). This crossover marks the transition from a coherently driven regime to an overdamped regime, in which stochastic detuning fluctuations dominate over coherent Rabi cycling. Beyond this crossover, reliable coherent control—such as the implementation of well-defined $\pi/2$ pulses within a single coherence time—becomes impractical due to rapid noise-induced dephasing. For the mechanically decoupled class of hBN emitters studied in Ref. [1], this limit occurs at
@@ V Acknowledgment
Supplementary Information
Hybrid Jump–Diffusion Model for Coherent Optical Control in hBN Quantum Emitters
Saifian Farooq Bhat, Michael K. Koch, Sachin Negi, Alexander Kubanek, and Vibhav Bharadwaj
I Numerical Implementation of the Jump–Diffusion Model
The stochastic evolution of the emitter transition frequency is implemented using a discrete-time realization of the Ornstein–Uhlenbeck (OU) process augmented by a Gaussian random jump (GRJ) mechanism. The update equation is
@@ I Numerical Implementation of the Jump–Diffusion Model
The time step was fixed at $dt=$1\text{\times}{10}^{-3}\text{\,}\mathrm{ns}$\$ to ensure stable numerical integration. The simulation window for each temperature point was $10\text{\,}\mathrm{ns}$, sufficiently long to sample both the OU fluctuations and discrete jumps. The evolution was implemented using the Euler–Maruyama scheme. Each realization consisted of $10^{5}$ stochastic trajectories, and ensemble averaging was used to extract linewidths and photon-correlation functions.
II Independent Parameter Study: Ornstein–Uhlenbeck Spectral Diffusion
To isolate the role of continuous spectral diffusion independently from jump processes, we first analyze the Ornstein–Uhlenbeck (OU) model as a baseline description of phonon-induced frequency noise. The OU process governing the instantaneous detuning $\omega(t)$ is written as

### v2

# A Hybrid Jump-Diffusion Model for Coherent Optical Control of Quantum Emitters in hBN

@@ Abstract
Hexagonal boron nitride (hBN) has emerged as a promising two-dimensional host for stable single-photon emission owing to its wide bandgap, high photostability, and compatibility with nanophotonic integration. We present a simulation-based study of temperature-dependent spectral dynamics and optical coherence in a mechanically decoupled quantum emitter in hBN. Employing a hybrid stochastic framework that combines Ornstein-Uhlenbeck detuning fluctuations with temperature-dependent, Gaussian-distributed discrete frequency jumps, motivated by experimentally observed spectral diffusion and blinking, we reproduce the measured evolution of inhomogeneous linewidth broadening and the progressive degradation of photon coherence across the relevant cryogenic range (5-30 K). The model captures phonon-related spectral diffusion with a cubic temperature dependence and the onset of jump-like spectral instabilities at higher temperatures. By calibrating the hybrid diffusion, jump parameters to the experimentally measured full width at half maximum (FWHM) of the emission line and analyzing the second-order correlation function $g^{(2)}(\tau)$ under resonant driving, we establish a unified phenomenological description that links stochastic detuning dynamics to the decay of optical coherence in a resonantly driven emitter. Analysis of $g^{(2)}(\tau)$ under resonant driving reveals an additional dephasing rate $\gamma_{\mathrm{sd+j}}$ that rises monotonically with temperature and drive strength, leading to a predicted critical crossover to overdamped dynamics at $T_{\mathrm{crit}}\approx 25.91$ K. This hybrid framework provides a quantitative connection between accessible spectroscopic observables and the dominant noise mechanisms limiting coherent optical control in mechanically decoupled quantum emitters, exemplified in hBN and generalizable to similar emitters in other materials.
@@ I Introduction
hBN possesses a wide bandgap of approximately $5.9~\text{eV}$ [1, 2], supporting deep-ultraviolet to near-infrared emission arising from defect-associated electronic transitions [3, 4, 5]. Owing to their brightness, polarization stability, and integrability, hBN quantum emitters have emerged as promising alternatives to semiconductor quantum dots for stable single-photon sources, provided that appropriate device engineering (encapsulation, substrate choice, or mechanical decoupling) is applied to mitigate spectral diffusion and sample variability [6, 7, 8]. In mechanically decoupled quantum emitters in hBN, reduced coupling to substrate-induced strain and charge noise has been shown to result in enhanced spectral stability and narrow optical resonances at cryogenic temperatures, including near 660 nm in representative system [9]. This mechanical isolation enables coherent optical control and exposes intrinsic phonon-limited coherence that is otherwise masked in substrate-bound samples. Despite these advantages, the optical coherence of hBN defect centers remains strongly influenced by phononic and electrostatic fluctuations. Temperature-dependent spectroscopy has shown that spectral diffusion, phonon coupling, and charge-induced dephasing can jointly broaden the emission linewidth and degrade coherent control of emitters in hBN [10, 11, 12]. For specific classes of emitters [9], both the zero-phonon line (ZPL) shift and linewidth have been reported to follow a cubic temperature dependence, consistent with deformation-potential coupling to low-energy acoustic phonons [13]. However, alternative temperature scalings have been reported for other hBN emitters, reflecting defect-specific phonon-coupling mechanisms and local environments [14]. Photoluminescence excitation spectroscopy further reveals that, in such emitters, an energy gap between the ZPL and the first acoustic mode of the phonon sideband can close with increasing temperature, leading to a rapid increase in dephasing once phonon protection is lost [9]. Beyond homogeneous phonon-induced broadening, hBN emitters also exhibit significant inhomogeneous broadening caused by slow spectral diffusion. Local charge reconfigurations, strain fluctuations, and defect rearrangements induce stochastic frequency shifts that broaden time-averaged spectra beyond the lifetime-limited linewidth, even at cryogenic temperatures. For the mechanically decoupled quantum emitter in hBN that serves as the experimental reference throughout this work (Ref. [9]), the measured linewidth follows
@@
with $A=1.01~\mathrm{GHz}$ and $B=3.77\times 10^{-5}~\mathrm{GHz/K}^{3}$. The constant term reflects radiative lifetime broadening together with residual quasi-static spectral diffusion, while the cubic term arises from acoustic-phonon-mediated dephasing. Mechanical decoupling suppresses the temperature-independent inhomogeneous component, but at higher temperatures, thermally activated charge motion and lattice fluctuations reintroduce spectral wandering and discrete frequency jumps that dominate the emitter’s dynamics (see Supplementary Information, Sec. III). Coherent optical control–including Rabi oscillations and $\pi$-pulse excitation–has been demonstrated up to $\sim 30~\mathrm{K}$ in mechanically isolated hBN [9, 15, 16]. At elevated temperatures, however, activated lattice noise and charge fluctuations lead to a transition from coherent to overdamped behavior governed by both continuous spectral wandering and abrupt jump-like instabilities [17, 18]. Reports of near-Fourier-limited coherence approaching room temperature in certain hBN-based emitter platforms, including suspended and nanophotonic device geometries, highlight the strong sensitivity of coherence preservation to fabrication quality, suspension geometry, and the local nano-environment [10, 19, 16, 9]. In this work, we develop a hybrid stochastic model inspired by recent theoretical treatments of spectral diffusion in solid-state emitters [20, 21], combining Ornstein-Uhlenbeck diffusion with a discrete Gaussian jump process to capture both continuous and abrupt detuning fluctuations (see Supplementary Information, Sec. I for numerical implementation and Sec. III for the physical motivation of jumps). Using experimentally extracted temperature-dependent parameters, the model reproduces the measured linewidth evolution, the power-dependent damping of optical Rabi oscillations, and the temperature-dependent behavior of the second-order correlation function $g^{(2)}(\tau)$ (see Supplementary Information, Fig. S2 and Sec. VII). This unified framework links phonon-driven diffusion with jump-like spectral instabilities and enables quantitative simulation of coherence dynamics in mechanically decoupled hBN emitters.
@@ II Theoretical Model and Simulation Framework
The spectral dynamics of a single quantum emitter in hBN are modeled using a hybrid stochastic approach that captures both continuous and discrete detuning fluctuations. This framework is motivated by recent theoretical treatments of spectral diffusion in solid-state emitters [20, 21] and by foundational studies of Gaussian and telegraph-like noise processes in semiconductor defects. The model combines Ornstein-Uhlenbeck (OU) diffusion, representing phonon-induced dephasing, with a Gaussian Random Jump (GRJ) process describing abrupt frequency shifts associated with charge motion, defect rearrangements, or local strain reconfigurations (see Supplementary Information, Sec. III). In the discrete form implemented numerically, the instantaneous detuning $\omega_{i}$ evolves according to
@@
where $dW_{t}$ is the Wiener increment and $J(t)$ is a compound Poisson process with rate $\lambda_{J}$ and Gaussian jump amplitudes. This hybrid diffusion-jump dynamics produces an effective inhomogeneously broadened emission line whose full width at half maximum (FWHM) is approximated by
@@
as measured in the mechanically decoupled hBN emitter of Koch et al. [9] (see Supplementary Information, Sec. VI and Table I for the complete two-stage calibration protocol using physically motivated temperature-dependent multipliers). The parameters $(S,\lambda_{J},\sigma_{J})$ are treated as temperature dependent and increase monotonically with $T$. For each temperature, Eq. (2) is integrated numerically using the Euler-Maruyama method over a 10 ns window, and the resulting detuning histograms are fitted to extract $\Gamma_{\mathrm{sim}}(T)$ (see Supplementary Information, Sec. V).
@@ III Simulation Results
The calibrated jump-diffusion model reproduces the experimentally observed temperature dependence of the emission linewidth of the specific mechanically decoupled hBN quantum emitter studied in Ref. [9]. For each temperature, the stochastic parameters $(S,\lambda_{J},\sigma_{J})$ were obtained through a constrained calibration in which their values were adjusted monotonically to match the experimentally measured inhomogeneous broadening $\Gamma_{\mathrm{exp}}(T)=A+BT^{3}$ with an accuracy better than $0.02~\mathrm{GHz}$ (see Supplementary Information, Sec. VI). The calibrated parameters provide qualitative insight into the relative contributions of continuous diffusion and jump-like detuning processes, and the resulting agreement between simulation and experiment is shown in Fig. 1, where the simulated linewidths reproduce the cubic, phonon-limited trend across the experimentally relevant temperature range from 5 to 30 K.
@@
Figure 2: (a-c) Simulated second-order correlation functions $g^{(2)}(\tau)$ for the hBN quantum emitter studied in Ref. [9] at 5 K, 20 K, and 30 K, respectively. Increasing excitation power increases the Rabi frequency ($\Omega_{R}\propto\sqrt{P}$), which manifests in $g^{(2)}(\tau)$ as more pronounced oscillatory shoulders and a higher oscillation frequency at short delay times. At elevated temperatures, these oscillatory features are progressively damped due to enhanced spectral diffusion and thermally activated frequency jumps, leading to suppression of coherent Rabi dynamics (see Supplementary Information, Fig. S2 for parameter dependence).
The effective decay rate extracted from the simulated photon-correlation traces provides a direct quantitative measure of how spectral diffusion and jump processes modify coherent optical dynamics. The additional dephasing term $\gamma_{\mathrm{sd+j}}$ appearing in Eq. (6) encapsulates the combined influence of continuous OU-type spectral diffusion and discrete Gaussian jumps–precisely the same mechanisms responsible for the inhomogeneous broadening discussed earlier in the linewidth analysis. As shown in Fig. 3, $\gamma_{\mathrm{sd+j}}$ is negligible at 5 K, resulting in decay rates that remain nearly independent of the Rabi frequency. At 20 K, $\gamma_{\mathrm{sd+j}}$ increases from a negligible value and develops a weak but measurable linear dependence on the Rabi frequency $\Omega_{R}$, coinciding with a fivefold increase in the calibrated jump amplitude $\sigma_{J}$, while the diffusion strength $S$ and jump rate $\lambda_{J}$ remain unchanged. At 30 K, $\gamma_{\mathrm{sd+j}}$ increases further as the jump-induced detuning remains strong and the diffusion strength $S$ rises by approximately 10%, leading to a steep increase in the decay rate that surpasses the $\Gamma=\Omega_{R}$ boundary. This marks the transition to an overdamped regime ($\Gamma=2\times\Omega_{R}$), in which noise-induced detuning dominates over coherent Rabi cycling. The resulting monotonic, temperature-driven enhancement of $\gamma_{\mathrm{sd+j}}$ establishes a direct link between inhomogeneous broadening, set by the combined action of spectral diffusion and jump processes. The observed suppression of coherent dynamics under resonant driving (see Supplementary Information, Fig. S3, which demonstrates that only the hybrid model with discrete jumps reproduces the experimentally observed compact, near-Gaussian detuning distributions at elevated temperatures).
@@ IV Conclusion
Our hybrid jump-diffusion model reproduces the experimentally observed temperature and power-dependent coherence degradation of the mechanically decoupled hBN quantum emitter studied in Ref. [9], using only the measured $T^{3}$ inhomogeneous broadening law as input [9]. The framework captures the interplay between continuous spectral diffusion and discrete jump-like detuning processes responsible for coherence loss under resonant driving. While the numerical values of the calibrated parameters are specific to this emitter, the modeling approach can be extended to mechanically decoupled hBN quantum emitters exhibiting similar temperature-dependent inhomogeneous broadening. Full numerical implementation, parameter scans, calibration strategy, analytical derivations, and validation are provided in the Supplementary Information (Secs. I-VII, Eqs. (1)-(9), Figs. S1-S3, and Table I). Inversion of Calibration and Extraction of Critical Dephasing Temperature
Most importantly, the model identifies the fundamental crossover where the effective dephasing rate equals the Rabi frequency (slope = 1). This crossover marks the transition from a coherently driven regime to an overdamped regime, in which stochastic detuning fluctuations dominate over coherent Rabi cycling. Beyond this crossover, reliable coherent control–such as the implementation of well-defined $\pi/2$ pulses within a single coherence time–becomes impractical due to rapid noise-induced dephasing. For the mechanically decoupled class of hBN emitters studied in Ref. [9], this limit occurs at
@@ Acknowledgment
Supplementary Paper
Hybrid Jump-Diffusion Model for Coherent Optical Control in hBN Quantum Emitters
I NUMERICAL IMPLEMENTATION OF THE JUMP-DIFFUSION MODEL
The stochastic evolution of the emitter transition frequency is implemented using a discrete time realization of the Ornstein-Uhlenbeck (OU) process augmented by a Gaussian random jump (GRJ) mechanism. The update equation is
@@ I NUMERICAL IMPLEMENTATION OF THE JUMP-DIFFUSION MODEL
The time step was fixed at $dt=$1\text{\times}{10}^{-3}\text{\,}\mathrm{ns}$\$ to ensure stable numerical integration. The simulation window for each temperature point was $10\text{\,}\mathrm{ns}$, sufficiently long to sample both the OU fluctuations and discrete jumps. The evolution was implemented using the Euler-Maruyama scheme. Each realization consisted of $10^{5}$ stochastic trajectories, and ensemble averaging was used to extract linewidths and photon correlation functions.
II Independent Parameter Study: Ornstein-Uhlenbeck Spectral Diffusion
To isolate the role of continuous spectral diffusion independently from jump processes, we first analyze the Ornstein-Uhlenbeck (OU) model as a baseline description of phonon induced frequency noise. The OU process governing the instantaneous detuning $\omega(t)$ is written as